myininaya (myininaya):

i have a fun algebra problem(i cant tell you where i seen this problem) solve for a+b a^2+2ab+b^2+4a+4b+4=0

OpenStudy (anonymous):

a^2+2ab+b^2+4a+4b+4=0 (a+b)^2 + 4(a+b) +4 =0 is a quadratic in (a+b)

myininaya (myininaya):

lol gj estudier

myininaya (myininaya):

you are ready for the _____ test

OpenStudy (dumbcow):

i agree...(a+b) = -2

OpenStudy (anonymous):

a+b = -2

myininaya (myininaya):

this was a fun question i wasnt expecting it on the ___ test

OpenStudy (dumbcow):

you can't say what test

myininaya (myininaya):

lol it said i'm not suppose to reveal any questions that i see on the test